OpenStudy (anonymous):

The table below shows the prices and discounts for two items. Item P- Rm 24 (Discount 25%) Item Q- Rm 30 (Discount x%) The total price of both items P and Q is RM 44.40 The value x is ? A 8 B  12 C 15 D 18 with working please.

OpenStudy (anonymous):

24(100-25)%+30(100-x)%=44.40 24(75/100)+30[(100-x)/100]=44.40 18+[(3000-30x)/100]=44.40 3000-30x=(44.40-18)*100 3000-30x=2640 30x=3000-2640 30x=360 x=12

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

now u r @ what sams??

OpenStudy (anonymous):

i mean i ur college

OpenStudy (anonymous):

I m in form 1 much more younger than most of u here! If you wants to know me better dont ask here!

OpenStudy (anonymous):

u r malaysian right??

OpenStudy (anonymous):

errrr....... You're smart!

OpenStudy (anonymous):

i also a malaysian... but i think smaller than u all.. coz i m secondary school student...

OpenStudy (anonymous):

me too la... I also secondary. Which school ar u in?

OpenStudy (anonymous):

SMK Munshi Abdullah Melaka

OpenStudy (anonymous):

what abt u??

OpenStudy (anonymous):

ohh, melaka. My friend name is Malcom Neoh, thats why one girl thought u were him :)

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

can we like, dont comment here??? Talk somewhere else pls... everyone can see our conversation la

OpenStudy (anonymous):

i'm a boy...

OpenStudy (anonymous):

ok, i know

OpenStudy (anonymous):

then do you know that i m a girl??

OpenStudy (anonymous):

nvr mind b4 leave delete all what u think is important

OpenStudy (anonymous):

ok, i think none of it! Do you have a phone? call me

OpenStudy (anonymous):

i know from ur pic

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

call u??

OpenStudy (anonymous):

why cant? then i hav 1 new friends

OpenStudy (anonymous):

num??

OpenStudy (anonymous):

apa???

OpenStudy (anonymous):

what is that???

OpenStudy (anonymous):

no hp... larx

OpenStudy (anonymous):

okkk, what evar!! GUD NITE!

OpenStudy (anonymous):

.... ok Gd nite... fb contact neoh124578@hotmail.com but i will nt always online... coz hv no time an gt many function nwadays... want to off lerx

OpenStudy (anonymous):

what??? You post yr fon no here? later everyone call u, you become very busy later wor... Hehheehehhe

OpenStudy (anonymous):

i delet larx.....

OpenStudy (anonymous):

and u sure its yrs not ur mum or dad or grandparents??

OpenStudy (anonymous):

yaya... i ald receive ur msg...

OpenStudy (anonymous):

me??

OpenStudy (anonymous):

how u know is me? Cannot be other people one ah??

OpenStudy (anonymous):

yaya the 012xxxx132 right??

OpenStudy (anonymous):

why xxxxxxxx one??

OpenStudy (anonymous):

coz gt 1 ppl msg me nw... n ask me know him or her or nt...

OpenStudy (anonymous):

hehe, GHOST! not me!

OpenStudy (anonymous):

how i know u want to post ur num or nt...

OpenStudy (anonymous):

oo

OpenStudy (anonymous):

see... i ald delet my hp num ald...

OpenStudy (anonymous):

**delete

OpenStudy (anonymous):

sry ar.. nw my hp no crdt... i will recharge afterwards